A 96-inch board has to be cut into tow pieces, one piece is 8 inches longer than the other. How long should each piece be?
Answer by stanbon(75887)

A 96-inch board has to be cut into two pieces, one piece is 8 inches longer than the other. How long should each piece be?
-------
Equation:
a = b + 8
a + b = 96 inches
----
Substitute for "a" and solve for "b":
b+8 + b = 96
2b = 78
b = 39 inches
-----------------
Solve for "a":
a = b + 8
a = 47 inches
======================
Cheers,
Stan H.
======================